# RoPE-LIME: RoPE-Space Locality + Sparse-$K$ Sampling for Efficient LLM Attribution

### v1

@@ Abstract
Explaining closed-source LLM outputs is challenging because API access prevents gradient-based attribution, while perturbation methods are costly and noisy when they depend on regenerated text. We introduce RoPE-LIME, an open-source extension of gSMILE that decouples reasoning from explanation: given a fixed output from a closed model, a smaller open-source surrogate computes token-level attributions from probability-based objectives (negative log-likelihood and divergence targets) under input perturbations. RoPE-LIME incorporates (i) a locality kernel based on Relaxed Word Mover’s Distance computed in RoPE embedding space for stable similarity under masking, and (ii) Sparse-$K$ sampling, an efficient perturbation strategy that improves interaction coverage under limited budgets. Experiments on HotpotQA (sentence features) and a hand-labeled MMLU subset (word features) show that RoPE-LIME produces more informative attributions than leave-one-out sampling and improves over gSMILE while substantially reducing closed-model API calls.
@@ 3.4 RoPE-LIME Pipeline
Attributions are computed by solving a weighted linear regression,
$\hat{\bm{\beta}}=\arg\min_{\bm{\beta}}\left\|\mathbf{W}^{1/2}\big(\mathbf{y}^{\mathrm{reg}}-\mathbf{Z}\bm{\beta}\big)\right\|_{2}^{2},\qquad a_{i}=\frac{|\beta_{i}|}{\sum_{j}|\beta_{j}|},$
where $\mathbf{Z}$ denotes Sparse-$K$ perturbations and $\mathbf{W}$ enforces locality via RWMD. Full algorithmic details are provided in Appendix A.2.
@@ A.1 Sparse-K parameter sweep
A.2 RoPE-LIME pipeline
$y_{j}^{\mathrm{reg}}=\mathbb{KL}(\mathcal{L}_{0}||\mathcal{L}_{j})$
Attribution scores $a_{j}$ quantify the contribution of each feature $F_{j}$; the intercept $\beta_{0}$ is omitted as it has no feature level interpretation.

### v2

@@ Abstract
Explaining closed-source Large Language Model (LLM) outputs is challenging because API access prevents gradient-based attribution, while perturbation methods are costly and noisy when they depend on regenerated text. We introduce Rotary Positional Embedding Linear Local Interpretable Model-agnostic Explanations (RoPE-LIME), an open-source extension of gSMILE that decouples reasoning from explanation: given a fixed output from a closed model, a smaller open-source surrogate computes token-level attributions from probability-based objectives (negative log-likelihood and divergence targets) under input perturbations. RoPE-LIME incorporates (i) a locality kernel based on Relaxed Word Mover’s Distance computed in RoPE embedding space for stable similarity under masking, and (ii) Sparse-$K$ sampling, an efficient perturbation strategy that improves interaction coverage under limited budgets. Experiments on HotpotQA (sentence features) and a hand-labeled MMLU subset (word features) show that RoPE-LIME produces more informative attributions than leave-one-out sampling and improves over gSMILE while substantially reducing closed-model API calls.
@@ 3.4 RoPE-LIME Pipeline
$y_{j}^{\mathrm{reg}}=\mathbb{KL}(\mathcal{L}_{0}||\mathcal{L}_{j})$
Attribution scores $a_{j}$ quantify the contribution of each feature $F_{j}$; the intercept $\beta_{0}$ is omitted as it has no feature level interpretation..